The perimeter of an isosceles triangle is 70dm. If the length of a leg of the triangle is 5dm less than twice the length of the base, find the lengths of the sides of the triangle.: The perimeter of an isosceles triangle is 70dm. If the length of a leg of the triangle is 5dm less than twice the length of the base, find the lengths of the sides of the triangle.
Answer by edjones(2169)

A) 2L+b=70
B) L=2b-5
.
A)
2(2b-5)+b=70 Substitute 2b-5 for L.
4b-10+b=70
5b=80
b=16 dm
.
B)
L=2*16 -5
=27 dm
.
27+27+16=70
.
Ed